OpenStudy (anonymous):

find the vertex and line of symmetry of f(x)=x^2+6x+3

Directrix (directrix):

(-3,-6) line of symmetry x = -3

Directrix (directrix):

Questions?

OpenStudy (anonymous):

wrong answer

OpenStudy (anonymous):

it states the vertex is (3,12)

Directrix (directrix):

The x-coordinate of the vertex is -b/2a = -6/2 = -3. y would be (-3)^2 + 6(-3) + 3 = 9 -18 + 3 = -6 With respect to you, I see no reason why the answer I gave is wrong. And, the line of symmetry is determined by the x-coordinate of the vertex. That would be x = -3. You did not mention that the line of symmetry is incorrect.